find the equation of a line parallel to 5x+10y=3 through (0,1)
Answer by TimothyLamb(4379)

---
given:
parallel to: 5x+10y=3
through: (0,1)
---
5x + 10y = 3
10y = -5x + 3
y = (-1/2)x + 3/10
---
parallel slope = (-1/2) = -0.5
---
to solve the linear equation, copy and paste this (the given point and parallel slope):
0,1,-0.5
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (2, 0)
y-intercept = (0, 1)
slope = -1/2 = -0.5
---
answer:
the slope-intercept form of our new parallel line:
y = (-1/2)x + 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php